Q) What is the kinetic energy, in joules, of a molecule of oxygen moving at a speed of 255 m/s(use the appropriate prefix to avoid exponential notation)

a) The equation for kinetic energy is ½mv2, and is in units of joules. What are the component units for joules in terms of mass and velocity?

kg m^2/s^2

b) From the previous part, we know that the mass needs to be in units of kg. What is the mass of an oxygen molecule in kg?

mass= 5.31E-26 kg

c) Finally, using the mass you calculated in the previous part and the velocity given in the problem, calculate the kinetic energy of one oxygen molecule moving at 255 m/s. Note the units given: 1 aJ = 10-18 J.
Answer #1


K.E = 1/2 mv^2

S.I unit of energy is Joules

So unit of mass and velocity should be S.I

unit of mass = kg
unit of velocity = m/s
************
b)
molar mass of O2 = 32 g/mol

1 mol has 6.022 x10^23 molecules
so mass of 1 molecule O2 = [ 1 mol/6.022 x10^23 ]* 32 g/mol
= 5.31 x10^-23 g
=5.31 x10^-26 kg
***********************
c)

K.E = 1/2 [5.31 x10^-26 kg]*[255 m/s]^2 = 172641. x10^-26 J

Answer

K.E = 1.73E-21 J

= 1.73E-3 aJ
******************************